OpenStudy (anonymous):

solve i22 22 is an exponent

OpenStudy (lgbasallote):

remember what i said earlier? if the exponent is even (except 2) then it's automatically 1

OpenStudy (anonymous):

i kno but thts not one of my answer choices

OpenStudy (asnaseer):

@janessia this question also involves complex numbers - you really need to learn about them first before attempting questions on them

OpenStudy (anonymous):

my choices are 5, i3 (3 is a exponent), and 21+i

OpenStudy (lgbasallote):

\[\huge i^2 = -1\] \[\huge i^{odd} = -i\] \[\huge i^{even} = 1\]

OpenStudy (lgbasallote):

there must be something missing here

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

maybe i got the methods mixed up

OpenStudy (lgbasallote):

i hate these formulae

OpenStudy (anonymous):

@asnaseer uhm this is how i learn if u dnt wanna help me then fine but @lgbasallote is helping me so thanks

OpenStudy (lgbasallote):

ohh

OpenStudy (lgbasallote):

well i'll just do it step-by-step \[i^{22} = (i^2)^{11} = (-1)^{11}\]

OpenStudy (asnaseer):

@janessia I never said I won't help you. your last question involved complex numbers and you indicated you knew nothing about them. so I advised you learn a bit about them first and gave you a link to a site where you can learn more about them. sorry if you feel I wasn't helping.

OpenStudy (anonymous):

take the integer remainder when you divide 22 by 4 the pattern is this \[i^0=1\] \[i^1=i\] \[i^2=-1\] \[i^3=-i\] \[i^4=1\] and it repeats every 4

OpenStudy (lgbasallote):

yeah that formula was wrong sorry

OpenStudy (anonymous):

so for example \(i^{201}=i^1=1\) because 4 goes in to 200 evenly leaving a remainder of 1

OpenStudy (anonymous):

@asnaseer i do better learning on here i cant do it on my own its too confusing for me plus i have like 5 min to kno how to do this its a timed quiz

OpenStudy (anonymous):

sorry \(i^{201}=i^1=i\) is what i meant

OpenStudy (asnaseer):

that isn't really learning...

OpenStudy (lgbasallote):

ahh so it's \[\LARGE i^{22(mod 4)}?\] lol

OpenStudy (lgbasallote):

@satellite73

OpenStudy (anonymous):

it is for me how u gone tell em wats learning for me??? if thts how i learn @asnaseer then thats juss how i learn! so bye! get off my question so i can understand what they are talking about!!!!

OpenStudy (anonymous):

yes, if you want to be fancy

OpenStudy (asnaseer):

bye...

OpenStudy (lgbasallote):

just refer to satellite's explanation there

OpenStudy (anonymous):

can one of yall explain to me wat ur doing since that person had me disracted?

OpenStudy (anonymous):

distracted i meant

OpenStudy (lgbasallote):

first identify your exponent

OpenStudy (lgbasallote):

what is the exponent of i^22

OpenStudy (anonymous):

22

OpenStudy (anonymous):

four goes in to twenty two 5 times with a remainder of 2 therefore \(i^{22}=i^2=-1\)

OpenStudy (anonymous):

ohk i understand thnx